Circular No. 6103

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1994ad IN ESO 152-G26 R. H. McNaught, Anglo-Australian Observatory, reports his discovery

of an apparent supernova (mag ∼ 18) in ESO 152-G26 on an I plate taken on Nov. 10 by M. Hartley with the U.K. Schmidt telescope. Confirmation was obtained on Nov. 11 via a CCD image taken by D. I. Steel and G. J. Garradd with the 1.0-m reflector at Siding Spring Observatory. Coordinates for the supernova were measured by McNaught from the CCD image: α = 1^{h}47^{m}39^{s}.57, δ = −56^{o}17′30′′.0 (equinox 1950.0, uncertainty in each coordinate 0′′.4); SN 1994ad is situated in a spiral arm and offset 0′′.6 west and 33′′.9 north from the galaxy’s center. No image appears in this position on the SERC J or the European Southern Observatory B and R surveys. A nearby star (∼ 0.5 mag fainter than the supernova) has position end figures 42^{s}.55, 18′14′′.8.

GRO J0422+32 J. Orosz and C. Bailyn, Department of Astronomy, Yale University,

report: “Optical observations of the black-hole candidate GRO J0422+32 in quiescence confirm an orbital period of 5.06 hr and strongly suggest a velocity semi-amplitude for the secondary star near 400 km/s. This would imply a mass function for the primary star of ∼ 1.4 M_{⊙}. Photometry with the Kitt Peak 2.1-m telescope (+ Tek CCD) on Oct. 27–28 confirms the quiescent magnitudes reported on IAUC 6072 and reveals a double-humped ellipsoidal lightcurve with an amplitude of 0.1 mag in I and a period of 0.2107 ± 0.0012 day. Spectra obtained with the Kitt Peak 4-m reflector (+ RC spectrograph) on Nov. 5–7 show double-peaked Hα and Hβ emission lines superposed on a continuum displaying the broad TiO absorption bands typical of early-M-type stars. Profile fitting to the emission lines implies an outer-disk velocity of 460 ± 20 km/s; in previous x-ray novae, this value has proved to be 10–20 percent greater than the velocity semi-amplitude of the secondary. While velocity determinations could not be obtained for individual spectra, sums of twenty-nine 1800-s exposures (with velocity shifts determined from sinusoids with the period given above, zero phase at Nov. 5.0025 UT, and semi-amplitudes between 370 and 430 km/s) provided strong cross-correlations with a template star of spectral type M0 V.”

1994 November 11 Daniel W. E. Green